A sorority set up a charity to help raise money for a local children's support group. They sold chicken dinners for $8 and steak dinners for $10. if 142 people ate dinner at the fundraiser and the sorority raised $1326, how many people ordered chicken and how many people ordered steak?

Answer by FrankM(1040)

8C+10S=1326
C+S=142
10C+10S=1420
8C+10S=1326
2C=94
C=47
S=95